OpenStudy (anonymous):

How do I simplify and state any restrictions on the variable: (x/x^2+5x+6) - (2/x^2+3x+2)? (Please show me the steps)

OpenStudy (saifoo.khan):

factor both the denominators for it's LCM.

OpenStudy (anonymous):

restriction: x^2 + 5x + 6 can't be zero. you have to solve this to obtain values of x and same way with x^2 +32 + 2

OpenStudy (saifoo.khan):

\[\frac{x}{(x+2) (x+3)}-\frac{2}{(x+1) (x+2)}\]

OpenStudy (saifoo.khan):

\[LCM = (x+1)(x+2)(x+3)\]

OpenStudy (saifoo.khan):

can u solve it further?

OpenStudy (anonymous):

I think no.